OpenStudy (anonymous):

a $6,300.00 principal earns 6% interest compounded annually. After 3 years, what is the balance in the account. Please help get the answer and explain.

OpenStudy (zzr0ck3r):

A = (P+r/n)^(nt) for you n = 12 P = 6300 r = .06 t = 3

OpenStudy (anonymous):

A= P ( 1 + r/n) ^(nt) P = principal amount r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year

OpenStudy (zzr0ck3r):

I had to google to remember the exact formula as well.... lol we got the same page

OpenStudy (anonymous):

took it in my last math class actually haha but yeah i forgot the formula too

OpenStudy (anonymous):

Lmao, thanks.